# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v2

@@ C Empirical Investigation of $P(z)$
Firstly, in Figure 15(a), we revisit the examples from Figures 6(a) and 6(b), coloring each sample based on the value of $P(z)$. As depicted in Figure 15(a), $P(z)$ proves to be an efficient metric to separate in-distribution samples from out-of-distribution samples; those in-distribution have high values for $P(z)$, whereas those out-of-distribution exhibit values of $P(z)$ near zero.
@@
We further visualize the entire optimization trajectory for the examples in Figures 6(c) and 6(d), i.e., optimizing the whiteness reward for Stable Diffusion v1.5 with the prompt "black duck" in Figure 16. Specifically, from Figure 16 we can clearly see that the value of $P(z)$ gradually decreases, and the generated image also gradually diverges from the distribution associated with a black duck. Notably, at around $20$ steps, the value of $P(z)$ becomes near-zero, and at the same time, the generated image more closely resembles a blue duck rather than the specified black duck.
@@
Figure 16: Trajectory of $P(z)$ on optimizing whiteness reward.
@@ D.1 Algorithm Implementation
Here, we describe an efficient method to implement our proposed hybrid gradient estimators detailed in Section 4, along with the optimization process, by utilizing the built-in auto-differentiation in PyTorch [Paszke et al., 2019]. Specifically, suppose we wish to use $q$ samples to estimate the gradient in Equation (20); that is, we draw $q$ noise vectors for perturbation: $\xi_{1},...,\xi_{q}$. We then generate the corresponding samples $x_{i}=M_{\theta}(z+\mu\xi_{i})$ for $i=1,...,q$. At this point, we should compute the estimated gradient of the reward models in a non-differentiable mode as follows:
@@ D.2 Experiment Details
Details for Figure 1. In these examples, we only use the base model of SDXL [Podell et al., 2024] as the image diffusion model. We adopt the DDIM sampler with 50 steps and $\eta=1$ for generation, and optimize all the injected noise in the generation process, the same as most experiments in this work. The classifier-free guidance is set to 5.0. For each reward function, we adopt the same hyperparameters for the optimizer and regularization terms as the experiments in Section 5.1 and 5.2. Note that the "Image Lightness" and "Image Darkness" correspond to the whiteness reward and blackness reward used in 5.1, respectively. From top to bottom in Figure 1, the used prompts are listed as follows:
@@
Details for Section 5.1. In this experiment, to solve the PRNO problem as formulated in Equation (17), we employ the Adam optimizer [Kingma and Ba, 2014] with a learning rate of $0.01$. For optimization with regularization, we set the regularization coefficient $\gamma$ to $1$. To compute the minibatch stochastic gradient for the regularization term in Equation (17), we set the batch size $b$—the number of random permutations drawn at each step—to $100$. For each optimization run, we utilize a single A800 GPU, with the total memory consumption being approximately 15 GB.
Details for Section 5.2. In this second set of experiments, we continue using the Adam optimizer with a learning rate of $0.01$. For optimization with regularization, though, we reduce the regularization coefficient to $\gamma=0.1$ because optimizing these human-like reward functions is less susceptible to the OOD reward-hacking issue, while maintaining the batch size for the permutation matrix $b$ at 100. Each optimization run also uses a single A800 GPU, but the total memory consumption is around 20 GB. Regarding the baselines in Table 1, we implemented LGD [Song et al., 2023a] ourselves, following the algorithm from their paper on these reward models. For other baselines, we reuse the statistics presented in their corresponding papers.
@@ E Hyperparameters Analysis
To examine the effects of $k$, $b$, and $\gamma$ on mitigating the OOD (Out-Of-Distribution) reward-hacking problem, we revisit the experiment of optimizing blackness reward with the prompt "white <animals>" from Section 5.1. In Figure 18, we illustrate how these three hyperparameters influence both the reward and the consistency score (CS), across four different values.

### v3

@@ C Empirical Investigation of $P(z)$
Firstly, in Figure 14(a), we revisit the examples from Figures 5(a) and 5(b), coloring each sample based on the value of $P(z)$. As depicted in Figure 14(a), $P(z)$ proves to be an efficient metric to separate in-distribution samples from out-of-distribution samples; those in-distribution have high values for $P(z)$, whereas those out-of-distribution exhibit values of $P(z)$ near zero.
@@
We further visualize the entire optimization trajectory for the examples in Figures 5(c) and 5(d), i.e., optimizing the brightness reward for Stable Diffusion v1.5 with the prompt "black duck" in Figure 15. Specifically, from Figure 15 we can clearly see that the value of $P(z)$ gradually decreases, and the generated image also gradually diverges from the distribution associated with a black duck. Notably, at around $20$ steps, the value of $P(z)$ becomes near-zero, and at the same time, the generated image more closely resembles a blue duck rather than the specified black duck.
@@
Figure 15: Trajectory of $P(z)$ on optimizing brightness reward.
@@ D.1 Algorithm Implementation
Here, we describe an efficient method to implement our proposed hybrid gradient estimators detailed in Section 4, along with the optimization process, by utilizing the built-in auto-differentiation in PyTorch [Paszke et al., 2019]. Specifically, suppose we wish to use $q$ samples to estimate the gradient in Equation (20); that is, we draw $q$ noise vectors for perturbation: $\xi_{1},...,\xi_{q}$. We then generate the corresponding samples $x_{i}=M_{\theta}(z+\mu\xi_{i})$ for $i=1,...,q$. At this point, we should compute the estimated gradient of the reward functions in a non-differentiable mode as follows:
@@ D.2 Experiment Details
Details for Figure 1. In these examples, we only use the base model of SDXL [Podell et al., 2024] as the image diffusion model. We adopt the DDIM sampler with 50 steps and $\eta=1$ for generation, and optimize all the injected noise in the generation process, the same as most experiments in this work. The classifier-free guidance is set to 5.0. For each reward function, we adopt the same hyperparameters for the optimizer and regularization terms as the experiments in Section 5.1 and 5.2. From top to bottom in Figure 1, the used prompts are listed as follows:
@@
Details for Section 5.1. In this experiment, to solve the probability-regularized noise optimization problem as formulated in Equation (17), we employ the Adam optimizer [Kingma and Ba, 2014] with a learning rate of $0.01$. For optimization with regularization, we set the regularization coefficient $\gamma$ to $1$. To compute the minibatch stochastic gradient for the regularization term in Equation (17), we set the batch size $b$—the number of random permutations drawn at each step—to $100$. For each optimization run, we utilize a single A800 GPU, with the total memory consumption being approximately 15 GB.
Details for Section 5.2. In this second set of experiments, we continue using the Adam optimizer with a learning rate of $0.01$. For optimization with regularization, though, we reduce the regularization coefficient to $\gamma=0.1$ because optimizing these human-like reward functions is less susceptible to the OOD reward-hacking issue, while maintaining the batch size for the permutation matrix $b$ at 100. Each optimization run also uses a single A800 GPU, but the total memory consumption is around 20 GB. Regarding the baselines in Table 1, we implemented LGD [Song et al., 2023a] ourselves, following the algorithm from their paper on these reward functions. For other baselines, we reuse the statistics presented in their corresponding papers.
@@ E Hyperparameters Analysis
To examine the effects of $k$, $b$, and $\gamma$ on mitigating the OOD (Out-Of-Distribution) reward-hacking problem, we revisit the experiment of optimizing darkness reward with the prompt "white <animals>" from Section 5.1. In Figure 17, we illustrate how these three hyperparameters influence both the reward and the consistency score (CS), across four different values.